OpenStudy (anonymous):

For the reaction, NH3 (g) + HCl (g) → NH4Cl (s) run at 95°C, ΔS = -284 J/K, and ΔH = -176 kJ. What is ΔG for this reaction? -149,020 J -91,368 J -71, 488 J 104,336 J @Somy @Abmon98 @Abhisar

OpenStudy (somy):

whats S and whats G

OpenStudy (anonymous):

@Somy ΔGosystem = ΔHosystem - T ΔSosystem

OpenStudy (somy):

good girl you just answered your own question :3

OpenStudy (anonymous):

lol so i subtract right cuz when i did it gave me a different answer @Somy

OpenStudy (somy):

T - is temperature i suppose

OpenStudy (somy):

so did u multiply temp by S value?

OpenStudy (anonymous):

yes @Somy

OpenStudy (somy):

in Kelvin unit or Celsius?

OpenStudy (anonymous):

idk am not good at chem i told you @Somy

OpenStudy (somy):

do it in Kelvin unit Celsius + 273= Kelvin

OpenStudy (somy):

and only then do the calculation

OpenStudy (anonymous):

386 divided by 49984

OpenStudy (somy):

95+273= ....K \(ΔGosystem = ΔHosystem - T \times ΔSosystem\) \(ΔGosystem = -176 kJ - ...K \times -284 J/K\)

OpenStudy (somy):

you do know that multiplication and division is first only then its subtraction and addition - this is math

OpenStudy (anonymous):

104336 @somy answer D

OpenStudy (anonymous):

it was worng :( @Somy

OpenStudy (somy):

im not familiar with this this is the answer you get according to the formula you gave

OpenStudy (anonymous):

ik chem sucks @Somy

OpenStudy (abhisar):

@abb0t

OpenStudy (somy):

oh i got it

OpenStudy (somy):

First of all we are making mistake in unit ΔGosystem = ΔHosystem - T ΔSosystem ΔHosystem = kj T= K ΔSosystem= kj/K this is the thing the question gives unit of S as J/K which means we have to divide that value by 1000 only then we can get the right answer after u calculate it multiply the answer by 1000 = coz the answer will be in kj and options are in J

OpenStudy (somy):

ΔHosystem= -176 kJ T= 95+273= 368 K ΔSosystem= -284 J/K /1000= -0.284 kJ/K and thus ΔGosystem = ΔHosystem - T ΔSosystem ΔGosystem = -176 kJ - 368*(-0.284)

OpenStudy (abb0t):

?

OpenStudy (abhisar):

is it correct @abb0t ?

OpenStudy (abhisar):

OH ! yes it's correct

OpenStudy (somy):

@Abhisar ❤_❤ yup i just learnt it :3